OpenStudy (anonymous):

The midpoint of a line segment with end points as (2, 3) and (4, y2) is (3, 2). What is the value of y2? Answer -2 -1 1 7

OpenStudy (mathmate):

We know that the y-coordinate of the mid-term is the average of the other two, so 3+y^2=2 Can you solve for y?

OpenStudy (mathmate):

Sorry, it's (1/2)(3+y^2)=2

OpenStudy (anonymous):

no i dont know how

OpenStudy (mathmate):

(1/2)(3+y^2)=2 multiply both sides by two: (3+y^2)=2*2=4 subtract 3 from each side: y^2 = 4-3 =1